# Retrieving the hot CGM physics from the X-ray radial profile from eROSITA with an IlustrisTNG-based forward model

### v1

@@ 5 Summary
In this work, we forward model the measurements of the X-ray surface brightness profiles obtained by stacking at the optical galaxy positions of the LS DR9 photometric (Fullphot) galaxy catalogue, reported by Z24. We retrieve the contribution of the hot CGM from central galaxies from that of point sources and satellite galaxies. Our hot CGM forward model is based on TNG300 hydrodynamical simulations. The main results from this work are summarized as follows:
@@
    We test the impact of the underlying halo mass distribution on the TNG-based prediction for corresponding X-ray surface brightness profiles. We do so by generating multiple mock galaxy catalogues using LC-TNG300, matched in stellar mass and redshift to the Fullphot galaxy catalogue, however, with different underlying halo mass distributions (Sec. 3.3; Fig. 1). Namely, the three models obtained by varying the halo mass distribution are as follows: model $1$ leaves LC-TNG300 halo mass distribution unchanged, model $2$ and $3$ changes the underlying halo distribution by excluding the $10\%$ and $30\%$ most massive halos from the original LC-TNG300 halo mass distribution, respectively. We show that the shape and normalization of the hot CGM X-ray surface brightness profiles are impacted by varying the halo mass distributions, where discarding the most massive halos from the underlying halo distribution results in steeper profiles with lower normalization. More precisely, we find that a factor 2$\sim$ 2$\times$ 2 × increase in the mean value of the underlying halo mass distribution results in 4$\sim$ 4$\times$ 4 × increase in the X-ray luminosity from the hot CGM.
@@
    We fit for the stacked X-ray radial surface brightness profile by eROSITA around MW-mass galaxies from Zhang et al. (2024a) with our forward models. Our model contains two emitting components (Eq. 1-3): hot gas (around central galaxies and around satellite galaxies hosted by more massive halos) and X-ray point sources (X-ray binaries and Active Galactic Nuclei). For three forward models, we compute the X-ray luminosity from point sources, $L_{X,\ \rm PS}$, and CGM (see results in Tab. 1). Using the empirical estimates for the expected luminosity from XRB (Sec. 4.1) and AGN (Sec. 4.2) for MW-mass galaxies, we put constraints on the permissible values of $L_{X,\ \rm PS}$ contribution to the X-ray stack (see Fig. 1). This analysis favours model three, which has a mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$, implying that the hot CGM component allows for a point source component with a mean AGN luminosity that agrees with empirical estimates from the low redshift universe. We focus our results on model three for all the following discussions of the hot CGM.
@@
    By integrating the area under the mean X-ray surface brightness profile from the central galaxy hot CGM component within $R_{500c}$, we obtain an X-ray luminosity, $L_{X,\ \rm CGM}=1.69_{-0.91}^{+0.28}\times 10^{40}$ ergs/s. We also show the residual plot of the per cent deviation of the data from our model, where the discrepancies within the $50-105$ kpc range are within $15\%$ (Fig. 3). We find that at $\lesssim 40$ kpc, the hot CGM from central galaxies and the X-ray point sources emission from XRB and AGN each account for $40-50\%$ of the total X-ray emission budget, respectively (Fig. 5). At larger radii $>40$ kpc, the contribution from the emission around satellites dominates the total X-ray emission, thereby explaining the overall flattening in the measurements.
@@
    We compare our results with the other hot CGM measurements presented in Z24, based on a different optical galaxy catalogue, namely from the SDSS spectroscopic survey (see comparison in Fig. 6). Our TNG-based forward model of the hot CGM prediction broadly agrees with the hot CGM measurement from Z24. The $L_{X,\ \rm CGM}$ measured between the two works are consistent. We note that the halo mass distributions of the two samples are similar mean values, where the mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$ for our forward model $3$ and the mean $M_{\rm 200m}^{\rm CEN}\sim 3\times 10^{12}\ {\rm M}_{\odot}$. This result further emphasizes the importance of the underlying halo mass distribution when comparing hot CGM profiles across different observations and simulation-based models.
This work provides a novel technique to constrain the mean AGN X-ray luminosity of a galaxy sample jointly with the radial hot CGM gas distribution within the halo using the X-ray hot CGM (stacking) measurements as a new benchmark. Alongside the progress in our understanding of how various stellar and AGN feedback prescriptions impact the hot CGM’s properties (Lau et al. 2024; Medlock et al. 2025), here, we emphasize another vital ingredient when comparing simulations with X-ray observations: the sensitivity of the X-ray CGM properties to the underlying halo mass distribution, stellar mass and redshift. One of the outstanding challenges in the current paradigm of galaxy formation and evolution models implemented in hydrodynamical simulations is to jointly constrain the microscopic scales (e.g., subgrid model physics) and its impact on the diffuse gas within the halo (Crain & van de Voort 2023). Future work implementing the data-comparison strategy developed here on other state-of-the-art simulations, like EAGLE (Crain et al. 2015; Schaye et al. 2015), FLAMINGO (Schaye et al. 2023), Magneticum (Dolag et al. 2005; Beck et al. 2016), SIMBA (Davé et al. 2019), will provide observationally-motivated ranges on allowed X-ray AGN luminosity for the MW-mass scales. Comparing the AGN X-ray luminosity predictions retrieved from the methodology developed here (informed by hot CGM X-ray observations) with that predicted by the simulation itself will provide a new ground for recalibrating and improving the current landscape of sub-grid AGN modes (e.g., see Alexander & Hickox 2012 for a review). Additionally, future X-ray missions on the observation side, like Athena (Nandra et al. 2013), AXIS (Mushotzky et al. 2019), HUBS (Cui et al. 2020) will push our current detection limits to resolve the hot CGM at higher redshifts in X-rays. This would further our understanding of how observations compare to the spatially resolved hot gas distribution at MW-mass scales in simulations.

### v2

@@ 5 Summary
In this work, we have forward-modeled the measurements of the X-ray surface brightness profiles obtained by stacking at the optical galaxy positions of the LS DR9 photometric (Fullphot) galaxy catalog, reported by Z24. We retrieved the contribution of the hot CGM from central galaxies from that of point sources and satellite galaxies. Our hot CGM forward model is based on TNG300 hydrodynamical simulations. The main results from this work are summarized as follows:
@@
    We tested the impact of the underlying halo mass distribution on the TNG-based prediction for corresponding X-ray surface brightness profiles. We did so by generating multiple mock galaxy catalogs using LC-TNG300, matched in stellar mass and redshift to the Fullphot galaxy catalog; however, with different underlying halo mass distributions (Sec. 3.3; Fig. 1). Namely, the three models obtained by varying the halo mass distribution are as follows: model $1$ leaves the LC-TNG300 halo mass distribution unchanged, whereas models $2$ and $3$ change the underlying halo distribution by excluding the $10\%$ and $30\%$ most massive halos from the original LC-TNG300 halo mass distribution, respectively. We show that the shape and normalization of the hot CGM X-ray surface brightness profiles are impacted by varying the halo mass distributions, whereby discarding the most massive halos from the underlying halo distribution results in steeper profiles with lower normalization. More precisely, we find that a factor $\sim 2\times$ increase in the mean value of the underlying halo mass distribution results in an $\sim 4\times$ increase in the X-ray luminosity from the hot CGM.
@@
    We fit the stacked X-ray radial surface brightness profile by eROSITA around MW-mass galaxies from Zhang et al. (2024a) with our forward models. Our model contains two emitting components (Eq. 1-3): hot gas (around central galaxies and around satellite galaxies hosted by more massive halos) and X-ray point sources (XRBs and AGNs). For three forward models, we computed the X-ray luminosity from point sources, $L_{X,\ \rm PS}$, and CGM (see results in Tab. 1). Using the empirical estimates for the expected luminosity from XRBs (Sec. 4.1) and AGNs (Sec. 4.2) for MW-mass galaxies, we put constraints on the permissible values of the $L_{X,\ \rm PS}$ contribution to the X-ray stack (see Fig. 1). This analysis favors model three, which has a mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$, implying that the hot CGM component allows for a point source component with a mean AGN luminosity that agrees with empirical estimates from the low-redshift Universe. We focus our results on model three for all the following discussions of the hot CGM.
@@
    By integrating the area under the mean X-ray surface brightness profile from the central galaxy hot CGM component within $R_{500c}$, we obtain an X-ray luminosity of $L_{X,\ \rm CGM}=1.69_{-0.91}^{+0.28}\times 10^{40}$ ergs/s. We also show the residual plot of the percentage deviation of the data from our model, where the discrepancies within the 5010550-105 kpc range are within $15\%$ (Fig. 3). We find that at $\lesssim 40$ kpc, the hot CGM from central galaxies and the X-ray point sources emission from XRBs and AGNs each account for $40-50\%$ of the total X-ray emission budget, respectively (Fig. 5). At larger radii >40>40 kpc, the contribution from the emission around satellites dominates the total X-ray emission, thereby explaining the overall flattening in the measurements.
@@
    We compared our results with the other hot CGM measurements presented in Z24, based on a different optical galaxy catalog, namely from the SDSS spectroscopic survey (see comparison in Fig. 6). Our TNG-based forward model of the hot CGM prediction broadly agrees with the hot CGM measurement from Z24. The $L_{X,\ \rm CGM}$ measured between the two works is consistent. We note that the halo mass distributions of the two samples are similar mean values, where the mean $M_{\rm 200m}=3.5\times 10^{12}\ {\rm M}_{\odot}$ for our forward model $3$ and the mean $M_{\rm 200m}^{\rm CEN}\sim 3\times 10^{12}\ {\rm M}_{\odot}$. This result further emphasizes the importance of the underlying halo mass distribution when comparing hot CGM profiles across different observations and simulation-based models.
This work provides a novel technique to constrain the mean AGN X-ray luminosity of a galaxy sample jointly with the radial hot CGM gas distribution within the halo using the X-ray hot CGM (stacking) measurements as a new benchmark. Alongside the progress in our understanding of how various stellar and AGN feedback prescriptions impact the hot CGM’s properties (Lau et al. 2025; Medlock et al. 2025), here we emphasize another vital ingredient when comparing simulations with X-ray observations: the sensitivity of the X-ray CGM properties to the underlying halo mass distribution, stellar mass, and redshift. One of the outstanding challenges in the current paradigm of galaxy formation and evolution models implemented in hydrodynamical simulations is jointly constraining the microscopic scales (e.g., subgrid model physics) and their impact on the diffuse gas within the halo (Crain and van de Voort 2023). Future work implementing the data-comparison strategy developed here on other state-of-the-art simulations, such as EAGLE (Crain et al. 2015; Schaye et al. 2015), FLAMINGO (Schaye et al. 2023), Magneticum (Dolag et al. 2005; Beck et al. 2016), and SIMBA (Davé et al. 2019), will provide observationally motivated ranges on the allowed X-ray AGN luminosity for the MW-mass scales. Comparing the AGN X-ray luminosity predictions retrieved from the methodology developed here (informed by hot CGM X-ray observations) with those predicted by the simulation itself will provide a new ground for recalibrating and improving the current landscape of sub-grid AGN modes (see e.g., Alexander and Hickox 2012 for a review). Additionally, future X-ray missions on the observation side, such as Athena (Nandra et al. 2013), AXIS (Mushotzky et al. 2019), and HUBS (Cui et al. 2020), will push our current detection limits to resolve the hot CGM at higher redshifts in X-rays. This would further our understanding of how observations compare to the spatially resolved hot gas distribution at MW-mass scales in simulations.